OpenStudy (anonymous):

how do you solve the square root of 2x minus the square root of x+1 equals 1?

OpenStudy (anonymous):

\[\sqrt{2x}-\sqrt{x+1}=1\]?

OpenStudy (anonymous):

yes that's the problem

OpenStudy (anonymous):

you are going to have to square twice i would start with \[\sqrt{2x}=1+\sqrt{x+1}\] and then square both sides

OpenStudy (anonymous):

\[(\sqrt{2x})^2=(1+\sqrt{x+1})^2\]\[2x=1+2\sqrt{x+1}+x+1\]

OpenStudy (anonymous):

then would you combine like terms?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that gives \[2x=2+x+2\sqrt{x+1}\] then subtract \(x\) and \(2\) from both sides

OpenStudy (anonymous):

you get \[x-2=2\sqrt{x+1}\] and then you have to square again

OpenStudy (anonymous):

i hate radicals lol

OpenStudy (anonymous):

yeah it is a pain when there are two, because you have to square twice

OpenStudy (anonymous):

next time you get \[x^2-4x+4=4(x+1)\]

OpenStudy (anonymous):

now you have a quadratic equation to solve once you solve it, make sure to check your answers in the original equation, because they might not both work

OpenStudy (anonymous):

would you take 4(x+1) to the other side

OpenStudy (anonymous):

i would multiply out first, make it easier \[x^2-4x+4=4x+4\]

OpenStudy (anonymous):

then subtract from both sides, you get an easy equation to solve

OpenStudy (anonymous):

how would you solve x squared -8x? would it be all that plus 1 or something?

OpenStudy (anonymous):

factor and get \[x^2=8x=0\] \[x(x-8)=0\]

OpenStudy (anonymous):

so \(x=0\) or \(x=8\) don't forget to check them

OpenStudy (anonymous):

o ok i get it now and thanks so much!!

OpenStudy (anonymous):

yw